OpenStudy (anonymous):

solve using the zero principle of zero products (9t +5)(4t -12)=0

OpenStudy (anonymous):

t=-5/9,3

OpenStudy (anonymous):

(9t+5)(4t-12)=0 If any individual factor on the left-hand side of the equation is equal to 0, the entire expression will be equal to 0. (9t+5)=0_(4t-12)=0 Set the first factor equal to 0 and solve. (9t+5)=0 Remove the parentheses around the expression 9t+5. 9t+5=0 Since 5 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 5 from both sides. 9t=-5 Divide each term in the equation by 9. (9t)/(9)=-(5)/(9) Simplify the left-hand side of the equation by canceling the common factors. t=-(5)/(9) Set the next factor equal to 0 and solve. (4t-12)=0 Factor out the GCF of 4 from each term in the polynomial. (4(t)+4(-3))=0 Factor out the GCF of 4 from 4t-12. (4(t-3))=0 Divide both sides of the equation by 4. Dividing 0 by any non-zero number is 0. (t-3)=0 Set each of the factors of the left-hand side of the equation equal to 0. t-3=0 Since -3 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3 to both sides. t=3 The final solution is all the values that make (9t+5)(4t-12)=0 true. t=-(5)/(9),3